OpenStudy (thedj4jc):

Jimmy has seven times as many dimes as nickels. Their value is $3.75. How many nickels does he have? 35 15 5 x(.10) = dimes? 7x(.05) = nickels?

OpenStudy (anonymous):

15

OpenStudy (briellshaw):

3.75/7=? @sstylesprettygirl please dont give out the exact answers without an explanation to back it up

OpenStudy (thedj4jc):

0.54 is the simplified answer I got from that.

OpenStudy (texaschic101):

0.10d + 0.05n = 3.75 -- multiply this by 100 to get rid of decimals 10d + 5n = 375 d = 7n 10(7n) + 5n = 375 70n + 5n = 375 75n = 375 n = 5 d = 7n d = 7(5) d = 35 there is 35 dimes (which equals 3.50) and 5 nickels (which equals 0.25) for a total of 3.75

OpenStudy (briellshaw):

A is the correct answer

OpenStudy (thedj4jc):

Thank you!!

OpenStudy (texaschic101):

there are 5 nickels and 35 dimes....

OpenStudy (anonymous):

look @BriellShaw I was incorrect by mistake why do you feel the need to correct me? but thax for the info

OpenStudy (briellshaw):

Im sorry for the inconveince

OpenStudy (anonymous):

no it not a problem i was being serious thx fror the info i didnt know that we needed to give a reason for the answer @BriellShaw